Circular No. 7647

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVALIKE VARIABLE IN SCORPIUS K. Haseda, Toyohasi, Aichi, Japan, reports his discovery of a nova-

like variable star on T-Max 400 films taken with a 0.10-m f/4 twin patrol camera. K. Kadota, Ageo, Saitama, provides the following position for the variable from Haseda’s discovery photograph: α = 17^{h}57^{m}06^{s}.98, δ = −32^{o}23′05′′.0 (equinox 2000.0). H. Yamaoka, Kyushu University, reports that Digital Sky Survey images show a dim star near this position with red mag ∼ 17.5 and blue mag ∼ 20. Haseda’s patrol films indicate that the star was fainter than mag 12.0–12.9 (not seen) on 27 dates spanning 1996 Mar. 26–2001 Apr. 22.744 UT; it was detected on his films at the following magnitudes during the past month: May 13.617, 10.5; 17.664, 11.0; 28.624, 10.5; June 15.661, 10.4. K. Takamizawa, Saku-machi, Nagano, Japan, reports the following prediscovery observations on his unfiltered T-Max 400 patrol films: Apr. 26.737, [12.9; May 12.643, 10.2; 16.698, 11.2; 20.692, 11.0; 25.660, 10.3. T. Kato, Kyoto University, adds that Kadota found no object brighter than mag 14 on a CCD image taken on 2000 Feb. 24.842, and that a search of the 2MASS and MSX5C catalogs has not revealed any likely counterpart.

W. Liller, Viña del Mar, Chile, reports that he measured a broadband V magnitude of 10.68 ± 0.03 on June 21.150 UT, using nearby GSC 7382837 (V = 9.87, B − V = +0.44). A low-dispersion objective-prism spectrogram recorded with the CCD shows a very intense red line, presumably Hα, that is ∼ 5 times the brightness of the surrounding continuum. Hβ appears to be weakly present, indicating that the star is substantially reddened.

COMET C/2001 K5 (LINEAR) Ephemeris extension to IAUC 7634 (elements MPEC 2001-M24):

2001 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 30 15^{h}08.^{m}26 −19^{o}25.^{′}1 5.527 6.249 131.^{o}7 7.^{o}0 14.2
July 10 15 05.99 −18 19.6 5.619 6.210 121.4 8.0 14.2
20 15 04.79 −17 20.3 5.728 6.170 111.3 8.8 14.2
30 15 04.70 −16 27.5 5.848 6.132 101.5 9.3 14.2
Aug. 9 15 05.69 −15 41.4 5.975 6.094  91.9 9.6 14.2
19 15 07.69 −15 01.6 6.102 6.056  82.6 9.5 14.2
29 15 10.64 −14 27.5 6.227 6.019  73.5 9.3 14.3
Sept. 8 15 14.47 −13 58.3 6.344 5.982  64.6 8.8 14.3
18 15 19.07 −13 33.0 6.451 5.947  55.9 8.0 14.3
2001 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green